# TabiBERT:A Large-Scale ModernBERT Foundation Model and Unified Benchmarking Framework for Turkish

### v1

@@ 5.3 Results
Academic $72.44$ $71.78$ YTU-Cosmos-BERT $0.66$
Retrieval $75.44$ $74.84$ BERTurk $0.60$
Code Retrieval $56.95$ $54.54$ BERTurk $2.41$
@@
5.3.1 Challenges and Adaptations
During our development and evaluation of TabiBERT, we encountered several challenges that required careful consideration and adaptation of our methodology.
Pre-training Throughput and Pre-tokenization
During our initial pre-training trials, we encountered extremely low throughput on 8×H100 SXM GPUs, achieving around only 3030% of the speed reported in the ModernBERT paper. We also ran training speed experiments on 8xA100SXM and training speed decreased only very slightly, which led us to believe that the bottleneck is the data loader and H100 GPUs are not utilized at their full potential. After consulting with the ModernBERT authors 444github.com/AnswerDotAI/ModernBERT/issues/236, we learned that the original authors applied pre-tokenization process to the training data to improve the throughput. This pre-tokenization step resolved the throughput issue, allowing us to achieve the training speeds reported in the ModernBERT paper. This experience highlights the critical role of thoroughly documented pre-training procedures in ensuring both reproducibility and scalability. With this lesson in mind, we mindfully documented the pre-training and fine-tuning processes for our own model to facilitate future research and practical adoption.
Evaluation Dataset Selection
During our evaluation, we initially selected the most commonly used supervised fine-tuning datasets in Turkish NLP research. However, after closer examination, we discovered that some of these datasets exhibited quality issues, including inconsistent annotations, ambiguous labeling criteria, or insufficient dataset size. To ensure reliable evaluation, we carried out a manual review of all candidate datasets, evaluating their annotation quality, label consistency, and task relevance. We selectively included only the higher-quality datasets in our evaluation suite, prioritizing those with clear annotation guidelines and reliable labels that are widely recognized in Turkish NLP research. This experience made evident the lack of standardized evaluation benchmarks in Turkish NLP research. To address this gap, we developed TabiBench, a comprehensive benchmark specifically designed for Turkish NLP, as further elaborated in Section 4.4.
Evaluation Methodology
For token classification tasks, including named entity recognition and part-of-speech tagging, we initially employed subword-level evaluation, which proved problematic. Different models tokenize the same word differently, with some tokenizers producing more subword tokens per word than others. This tokenization discrepancy made subword-level scores incomparable across models, as model tokenizers with higher token fertility would be able to be score higher by trivially predicting more fragmented subword tokens per word. To address this, we modified our evaluation methodology to use word-level evaluation instead, ensuring that predictions are properly aggregated at word boundaries before computing metrics. This approach provides a fair comparison across models regardless of their tokenization strategies.
These challenges highlight the importance of careful dataset selection, appropriate evaluation methodology design, and metric selection tailored to each task’s specific requirements.
6 Conclusion and Future Work
This work introduced TabiBERT and TabiBench. TabiBERT is a Turkish encoder model that leverages the state-of-the-art architectural innovations of the ModernBERT. It is pre-trained from scratch on a large, diverse corpus of native Turkish texts from the web, scientific, literary, academic, and parliamentary domains, complemented with limited bilingual, code, and math data. It achieves state-of-the-art results across a wide range of Turkish NLP tasks while being efficient in inference, with high speed and a small memory footprint. The gains from its 16 times longer context support compared to earlier models are demonstrated across multiple tasks without sacrificing performance on short-context tasks. The adoption of a modern tokenizer with explicit support for code, math, and structured text broadens the applicability of Turkish encoder models compared to prior ones.
TabiBench is a new comprehensive benchmark for Turkish NLP that encompasses a diverse set of tasks, including text classification, sequence labeling, question answering, and retrieval. It is designed to serve as a standard evaluation suite for Turkish language understanding models, with carefully curated high-quality datasets. Moreover, it covers a wide range of domains, including social media, product comments, treebanks, news, web search, academic and medical texts, and code. Every task in TabiBench includes standard train, validation, and test splits, along with clear evaluation metrics for consistent assessment.
Experiments conducted on TabiBench demonstrate the effectiveness of TabiBERT. It outperforms existing monolingual Turkish models across a broad range of tasks and establishes a new state of the art in overall benchmark performance. As a result of large-scale pre-training on huge, multilingual corpora and cross-lingual transfer, multilingual models such as mmBERT may achieve higher absolute scores. However, our results point out that a monolingual encoder whose design decisions were carefully made provides a strong trade-off between efficiency and performance on Turkish-centric applications. The results show that by utilizing modern encoder architectures, extended context lengths, and large-scale pre-training on large, high-quality, diverse and language-specific datasets, substantial gains are achieved for Turkish, known for being a morphologically rich and underrepresented language. Beyond absolute scores, this work provides a standardized, reproducible evaluation methodology for Turkish NLP, addressing long-standing issues of inconsistent dataset splits and fragmented evaluation practices.
We hope that TabiBERT and TabiBench will serve as valuable resources for the Turkish NLP community, fostering further research and development in this area.
There are several promising directions for extending this work. TabiBench already covers a wide range of tasks; however, it can be expanded to include more, such as math and long context tasks, enabling further evaluation of TabiBERT under these conditions. Furthermore, its long context support can be exploited in processing longer texts in tasks such as document classification and long-context question answering. Moreover, a general, state-of-the-art text embedding model can be mid-trained starting from TabiBERT to enhance performance and support longer contexts for retrieval tasks. Finally, pre-training a larger version of TabiBERT with a BERT-large configuration could be explored to potentially achieve even better results across tasks.
@@ 7 Limitations
Pre-training
Given our computational limitations, we were able to pre-train only the base model, comprising 149149M parameters, whereas the ModernBERT family also includes a larger 395395M parameter variant. In addition, our pre-training corpus of 86.5886.58B tokens is modest compared to large corpora such as FineWeb, consisting of 1818T tokens. Moreover, TabiBERT is pre-trained for 11T tokens, compared to ModernBERT and mmBERT which were pre-trained for 22T and 33T tokens respectively. As a result of limited Turkish corpora, TabiBERT was trained for $11.83$ epochs on the same data. It can be argued that this might limit its overall generalization and expose it to risks of overfitting or memorization. However, we hypothesize that the limited model size and consequently, capacity should mitigate these risks to great extent. Larger models and more diverse, expensive corpora could further enhance the linguistic capabilities and robustness of Turkish language models.
Fine-Tuning Experiments
Due to time and compute budget constraints, we were not able to conduct neither statistical significance testing, nor conduct fine-tuning experiments with multiple runs for each task using different random seeds, which prevents us from reporting more stable averages for downstream performance. Moreover, a portion of our benchmark consists of machine-translated datasets from English, specifically, all retrieval and code retrieval tasks, as well as three of four academic domain tasks. These translations, despite being carried out by a state-of-the-art LLM such as GPT-4.1, while useful for benchmarking and coverage, may not fully capture the intricacies of native Turkish.
Long Context Capabilities
While our model is capable of processing long input contexts, up to $8.192$ tokens, which is $16$ times of other major Turkish models, due to the lack of Turkish finetuning tasks with long-context, we cannot demonstrate the full advantage of this architectural feature at this time. Expanding available Turkish datasets to better reflect real-world, high-context use cases remains an important area for future work.
Data
@@
Due to licensing restrictions on some source materials (academic publications, books), we cannot release the complete pre-training corpus. However, we provide detailed statistics, composition breakdowns, and preprocessing methodology to enable replication with alternative corpora.
8 Ethics Statement
Web content carries the risk of harmful content including bias, toxicity, abuse, and obscenity. In removing such harmful content from the training data, we rely on the methods used by the providers of the web corpus. Despite the filtering and cleaning efforts, there is still a risk that some harmful content remains in the training data. Although TabiBERT is primarily intended for use in text understanding tasks, the base model may still be used to generate text with the masked language modeling objective. Therefore, it is possible for TabiBERT to be affected by such harmful content and produce biased or toxic outputs. This calls for continuous monitoring of this system to eliminate such occurrences.
Heterogeneous sources such as academic texts, books, and parliamentary records were included in the training data to ensure linguistic coverage and representativeness for Turkish. Nevertheless, users should be mindful of ethical considerations related to data provenance and domain-specific bias when applying the model.
TabiBERT and TabiBench are released as research artifacts to support the development and evaluation of Turkish NLP systems. We encourage responsible use, task-specific evaluation, and bias-aware analysis when deploying the model in downstream applications.
@@ Acknowledgments
We thank VNGRS [^5] https://vngrs.com/ for funding the computational resources used for model pre-training and fine-tuning. We also thank Lambda Labs [^6] https://lambda.ai/ for providing additional GPU resources. Finally, we thank Selman Baysan for his contributions to the implementation of the retrieval tasks.
8.1 AI Assistance
@@ A.1 Fine-tuning Results
The appendix presents detailed performance comparisons across all evaluation tasks. Each table compares TabiBERT against established monolingual Turkish models with the best-performing monolingual model highlighted in each task. We also include mmBERT as a reference point. While the very large-scale multilingual model mmBERT often achieves competitive or superior performance, direct comparison is complicated by substantial differences in model scale, training data volume, and cross-lingual training objectives.
@@ A.1.1 Text Classification
Model performance is evaluated across four text classification datasets: NewsCat (McemilgNewsCat), BilTweetNews (Toraman2017EventPrediction; Toraman2021EventRetrieval), GenderHateSpeech (Toraman2022HateSpeech; Sahinuc2023GenderBias), and ProductReviews (FthbrmnbyTurkishProductReviews). Table 6 presents these results using the macro F1 score. With the exception of TurkishBERTweet, the models achieve comparable results across the selected tasks. The YTU-Cosmos-BERT model yields the highest performance with a macro F1 score of $84.25$, followed by TabiBERT at $83.44$. Although TabiBERT does not attain the highest score on individual benchmarks, its is comparable with the top-performing models. Notably, mmBERT underperforms relative to the Turkish models with an average macro F1 score of $82.54$.
@@ A.1.2 Token Classification
Token classification performance is evaluated using four datasets: WikiNER (Altinok2023Turkish), WikiANN TR (Rahimi2019WikiANN), PosUDBOUN (UDTurkishBOUN), PosUDIMST (UDTurkishIMST). Table 7 presents the micro F1 scores. All Turkish models demonstrate high performance, with micro F1 scores exceeding $92$. The BERTurk model achieves the highest overall score of $93.67$ taking the leading position across all datasets. Consistent with the results observed in text classification tasks, TabiBERT maintains competitive performance levels.

### v3

@@ 5.3.1 Overall Performance
Information Retrieval $75.44$ $74.84$ BERTurk $0.60$
Academic Understanding $72.44$ $71.78$ ytu-cosmos-BERT $0.66$
Code Retrieval $56.95$ $54.54$ BERTurk $2.41$
@@
5.3.2 Performance by Task Category
Text understanding tasks
Text classification: Results converge among top models (except TurkishBERTweet). ytu-cosmos-BERT leads (84.25), followed closely by TabiBERT (83.44) and BERTurk (83.18). Margins under 1 point suggest approaching saturation.
Token classification: All Turkish models exceed 92. BERTurk (93.67), ytu-cosmos-BERT (93.60), and TabiBERT (93.42) cluster within 0.25 points, indicating strong performance across models with limited improvement headroom.
Semantic textual similarity: Top models (excluding TurkishBERTweet) score 84.59–85.33. BERTurk leads (85.33), TabiBERT follows (84.74), then ytu-cosmos-BERT (84.59).
Natural language inference: TabiBERT achieves the highest Turkish model score (84.51), marginally exceeding BERTurk (84.33) and ytu-cosmos-BERT (84.16).
These results demonstrate that TabiBERT’s long-context optimization preserves sentence-level semantic precision, narrow margins across tasks indicate no performance sacrifice for extended context capability.
Question answering
[…]
Academic understanding tasks
TabiBERT leads Turkish models (77.02), exceeding BERTurk (76.36) and ytu-cosmos-BERT (76.12). Particularly strong on Thesis-Abstract-Classification-11K (187-class task with specialized language and limited per-class data), where pretraining on scientific articles and theses provides a domain advantage, in addition to the impact of its long context support.
Information retrieval
TabiBERT achieves 75.44 NDCG@10 on TR-MTEB benchmarks, surpassing BERTurk (74.84). Modest gain reflects the benefits of an extended context for processing longer documents without truncation.
Code retrieval
TabiBERT substantially outperforms Turkish models: 56.95 NDCG@10 versus BERTurk (54.54, +2.41) and ytu-cosmos-BERT (53.80, +3.15). Improvements stem from the extended context for longer code snippets, exposure to code during pretraining, and the tokenizer’s explicit code syntax support. For retrieval tasks requiring document-level coherence and cross-modal understanding, these design choices provide measurable advantages.
5.3.3 Comparison with Multilingual mmBERT
mmBERT achieves 79.26, a 1.68-point advantage over TabiBERT. This reflects fundamental resource differences rather than architectural superiority, as both share ModernBERT architecture with different specialization strategies.
Resource differences
[…]
Tokenization: mmBERT’s tokenizer shows 41 per cent higher fertility on Turkish text, requiring 41 per cent more tokens for identical content. This increases computational cost and reduces effective context by  29 per cent (equivalent to a reduction from $8,192$ to  $5,800$ tokens of Turkish text). Finer tokenization may capture morphological nuances via character n-gram fragmentation but sacrifices efficiency and effective context length.
Training scale: mmBERT pretrained on 3T tokens versus TabiBERT’s 1T, threefold of the latter. Massive multilingual corpus (1800 languages) enables cross-lingual transfer, particularly from high-resource languages like English. Different philosophy: mmBERT leverages cross-lingual knowledge; TabiBERT focuses computational resources on Turkish-specific phenomena.
Practical trade-offs
Despite mmBERT’s higher scores, TabiBERT offers advantages for Turkish-focused applications: faster inference (due to lower tokenizer fertility),  51 per cent smaller memory footprint, and 29 per cent longer effective context for Turkish text. These efficiency benefits suit resource-constrained deployments or low-latency applications.
Our results demonstrate that targeted, language-specific development with careful architectural choices can achieve competitive performance with substantially lower resource requirements. Model choice depends on application needs: Turkish-focused tasks may benefit from TabiBERT’s efficiency and specialized coverage; multilingual applications or those requiring cross-lingual transfer may prefer mmBERT.
6 Discussion
Figure 2 visualizes performance patterns across TabiBench. This section interprets these results, discusses architectural trade-offs, and addresses methodological considerations that informed our evaluation design.
6.1 Interpreting Performance Patterns
Task saturation versus architectural advantages
Results reveal a clear pattern: narrow margins (~1 point) on text understanding tasks (classification, token classification, STS, NLI) suggest these benchmarks approach saturation for base-sized encoders. All top Turkish models converge around similar scores, indicating limited remaining headroom. Critically, TabiBERT’s competitive performance on these tasks demonstrates that long-context optimization does not compromise short-text understanding, a key validation of ModernBERT’s architectural design.
Conversely, substantial gains on question answering (+9.55), code retrieval (+2.41), and academic understanding (+0.66) reveal where extended context and specialized pretraining provide measurable advantages. These tasks stress capabilities that saturated benchmarks cannot assess: processing long passages, maintaining coherence across extended contexts, and cross-modal understanding (natural language + code).
Architectural innovations successfully transfer to Turkish
TabiBERT’s consistent performance across diverse tasks validates that ModernBERT’s innovations, RoPE positional embeddings, Flash Attention, and optimized layer configurations transfer effectively to Turkish. The model achieves competitive or superior results while providing efficiency benefits (faster inference, lower memory) over earlier BERT-based Turkish models. This demonstrates that recent encoder architecture advances are not English-specific but generalize to morphologically rich, agglutinative languages.
Monolingual specialization versus multilingual scale
[…]
Curse of multilinguality
Despite the now well-known phenomenon of the “curse of multilinguality” (Conneau2020XLM-R), it is observed that mmBERT, a multilingual model, outperforms TabiBERT on several tasks. This can be attributed to the gains obtained by the factors, explained in Section 5, that contribute to this phenomenon, outweighing the negative effects of multilinguality in this case.
6.2 Methodological Considerations and Design Decisions
Tokenization-agnostic evaluation for token classification
For named entity recognition and part-of-speech tagging, we adopted word-level rather than subword-level evaluation. This addresses a fundamental comparability issue: different tokenizers fragment words into varying numbers of subword tokens. A higher-fertility tokenizer would achieve artificially inflated subword-level scores by correctly predicting more fragmented tokens per word, even with identical word-level accuracy.
Word-level evaluation aggregates predictions at word boundaries before computing metrics, ensuring fair comparison regardless of tokenization strategy. While subword-level metrics may be appropriate for analyzing tokenizer-specific behavior, cross-model comparisons require tokenization-agnostic metrics. This principle is broadly relevant in multilingual NLP, where tokenization strategies vary substantially.
Pretraining infrastructure and reproducibility documentation
[…]
This experience underscores the importance of comprehensive documentation for reproducible large-scale training. Critical but “obvious” implementation decisions must be explicitly stated; otherwise, researchers attempting replication may encounter substantial inefficiencies or failed reproductions. We have documented our procedures in detail (Section 3) to facilitate future work.
@@ 7 Limitations
We acknowledge limitations constraining this work’s scope and indicating areas for future research.
7.1 Computational Resource Constraints
Model scale
We pretrained only base-sized models (149M parameters) due to budget limitations. ModernBERT includes a large variant (395M parameters). Scaling experiments with BERT-large configuration could yield stronger performance, particularly on tasks requiring greater capacity (complex reasoning, low-resource specialized domains).
Corpus size and training duration
Our pretraining corpus (86.58B tokens) is modest compared to large-scale corpora like FineWeb (18T tokens). TabiBERT was pretrained for 1T tokens versus ModernBERT (2T) and mmBERT (3T). Limited Turkish text availability necessitated 11.83 epochs over the same data, raising potential overfitting concerns. However, we hypothesize that base model capacity (149M parameters) substantially mitigates memorization risks; that is, the model likely lacks sufficient capacity to memorize the corpus. Larger, more diverse Turkish corpora would directly address this limitation.
7.2 Evaluation Constraints
Statistical significance
Due to computational constraints, we did not conduct significance testing or multiple fine-tuning runs with different random seeds. Reported scores represent single runs. Performance margins <1 point (common on saturated tasks) may or may not reach statistical significance. Formal hypothesis testing with multiple runs would be required for definitive relative performance claims on close-call tasks. We present results as indicative evidence rather than statistically confirmed superiority.
7.3 Translation-Based Datasets: Rationale and Limitations
Thirteen of the twenty-eight TabiBench datasets consist of machine-translated content (all information and code retrieval tasks plus three academic understanding tasks). Seven of them were translated by us using GPT-4.1; six (TR-MTEB) were translated by prior work (BaysanGungor2025TR-MTEB). We opted to rely on translations from a pragmatic stance. High-quality Turkish datasets for retrieval and specialized domains are scarce or non-existent, yet these tasks are critical for comprehensive encoder evaluation. For code retrieval, only natural language components are translated while code remains unchanged. As a result of using a state-of-the-art LLM for translation, we present a high quality translation that preserves original meaning and task relevance.
Machine translated content may have potential limitations, such as native linguistic nuances, idiomatic expressions or cultural context not fully captured. However, in our cases, all of the content is technical or formal in nature, minimizing such risks. Moreover, omitting such task categories with the aim of avoiding machine-translated content would severely limit the evaluation scope.
Long-context evaluation gap
[…]
7.4 Data Availability
@@
Due to licensing restrictions on portions of our pretraining corpus (academic publications, books from institutional repositories), we cannot release complete training data. This limits exact reproducibility. However, we provide detailed corpus statistics, composition breakdowns, and preprocessing methodology (Section 3), enabling replication with alternative corpora. Researchers can assemble comparable Turkish corpora using publicly available web data, parliamentary records, and permissively licensed scientific literature following our documented procedures.
8 Conclusions and Future Work
This work introduces two resources for Turkish NLP: TabiBERT, a ModernBERT-based encoder demonstrating that recent architectural advances transfer effectively to Turkish; and TabiBench, a standardized evaluation benchmark addressing reproducibility challenges in Turkish NLP research.
8.1 Key Contributions
TabiBERT
[…]
TabiBERT achieves 77.58 average on TabiBench, establishing state-of-the-art among monolingual Turkish models with particular advantages in question answering (+9.55), code retrieval (+2.41), and academic understanding tasks (+0.66). Competitive sentence-level performance that demonstrates long-context optimization does not sacrifice short-text understanding. Substantial efficiency benefits: faster inference, lower memory ( 51 per cent smaller than mmBERT), 29 per cent longer effective context than higher-fertility multilingual alternatives.
[…]
TabiBench
We developed a standardized benchmark addressing Turkish NLP evaluation infrastructure gaps: 28 curated datasets across 8 task categories with fixed splits and task-appropriate metrics. Manual quality review ensured annotation consistency and adequate size. Fair comparison protocol fine-tunes all models on TabiBench, attributing performance differences to architecture rather than dataset-specific optimization.
Released as open resource to establish evaluation best practices, enable reproducible cross-model comparisons, and provide a template for similar efforts in other lower-resource languages. Addresses long-standing reproducibility issues in Turkish NLP.
Methodological insights
Documentation of pretraining infrastructure challenges highlights the importance of comprehensive implementation detail reporting for reproducibility. Adoption of tokenization-agnostic metrics (word-level for token classification) addresses fundamental comparability issues relevant across multilingual NLP.
8.2 Broader Implications
For Turkish NLP: Provides both updated model architecture and evaluation infrastructure as foundations for future research.
For lower-resource languages: Methodologies, systematic quality review, language-specific optimization, transparent constraint reporting, and generalize to other languages lacking standardized evaluation frameworks.
For model development philosophy: Supports viability of targeted, resource-efficient development as an alternative/complement to large-scale multilingual approaches. Optimal strategy depends on application requirements (multilingual coverage vs. specialized efficiency) rather than universal superiority.
8.3 Future Directions
Long-context evaluation datasets
Most significant gap: absence of Turkish datasets specifically evaluating long-context understanding (long and multi-document QA, legislative analysis). Developing such benchmarks would enable comprehensive, extended-context assessment and drive progress on applications requiring long-range understanding.
Statistical validation
Comprehensive significance testing across benchmarks with multiple seeds would formalize performance comparisons. Robustness analysis under input perturbations (typos, morphological variations, domain shifts) would deepen understanding of capabilities and failure modes.
Scaling experiments
Pretraining a larger variant (BERT-large, 395M parameters) would reveal whether advantages scale proportionally with capacity. Extended training beyond 1T tokens on expanded corpora would determine whether additional training yields proportional gains or diminishing returns.
Domain-specific optimization
Fine-tuning or continued training for specialized domains (legal, medical, scientific) could reveal differential benefits for applications requiring domain expertise.
Embedding model development
Continued training from TabiBERT optimized for retrieval could provide Turkish NLP with a specialized embedding model potentially superior to multilingual alternatives.
Cross-lingual comparative analysis
Systematic comparison of Turkish-specific versus multilingual ModernBERT variants, controlling for training compute and corpus quality, would quantify monolingual specialization benefits versus cross-lingual transfer.
TabiBench evolution
Benchmark can expand to additional task types, such as mathematical understanding, and maintain versioned updates as new datasets become available. Living benchmark with periodic releases would support sustained progress measurement.
TabiBERT and TabiBench contribute to Turkish NLP by providing both an efficient, modern encoder model and a standardized evaluation benchmark. These resources address infrastructure gaps that have hindered progress, and we believe they will facilitate more rigorous and reproducible research in Turkish language technology.
@@ Acknowledgments
We thank VNGRS [^6] https://vngrs.com/ for funding the computational resources used for model pretraining and fine-tuning. We also thank Lambda Labs [^7] https://lambda.ai/ for providing additional GPU resources. Finally, we thank Selman Baysan for his contributions to the implementation of the retrieval tasks.
AI Assistance
@@
Declaration of Competing Interests
The authors declare the following financial interests/personal relationships which may be considered as potential competing interests: Melikşah Türker is a PhD student at Boğaziçi University and is employed by VNGRS, which provided the cloud GPU infrastructure for model training. Dr. Onur Güngör is a part-time faculty member at Boğaziçi University and is also employed by Udemy. While these entities may promote this research for institutional branding or professional visibility, the authors received no direct financial benefits from these activities. All training, evaluation, and manuscript preparation were conducted independently by the research lab to ensure transparency and reproducibility. The companies had no role in the study design or the reporting of results. All other authors declare no competing interests.
@@ A.1 Fine-tuning Results
The appendix presents detailed performance comparisons across all evaluation tasks. Each table compares TabiBERT against established monolingual Turkish models, with the best-performing monolingual model highlighted in each task. We also include mmBERT as a reference point. While the very large-scale multilingual model mmBERT often achieves competitive or superior performance, direct comparison is complicated by substantial differences in model scale, training data volume, and cross-lingual training objectives.
@@ A.1.1 Text Classification
Model performance is evaluated across four text classification datasets: NewsCat (McemilgNewsCat), BilTweetNews (Toraman2017EventPrediction; Toraman2021EventRetrieval), GenderHateSpeech (Toraman2022HateSpeech; Sahinuc2023GenderBias), and ProductReviews (FthbrmnbyTurkishProductReviews). Table 6 presents these results using the macro F1 score. With the exception of TurkishBERTweet, the models achieve comparable results across the selected tasks. The ytu-cosmos-BERT model yields the highest performance with a macro F1 score of $84.25$, followed by TabiBERT at $83.44$. Although TabiBERT does not attain the highest score on individual benchmarks, it is comparable with the top-performing models. Notably, mmBERT underperforms relative to the Turkish models with an average macro F1 score of $82.54$.
@@ A.1.2 Token Classification
Token classification performance is evaluated using four datasets: WikiNER (Altinok2023Turkish), WikiANN TR (Rahimi2019WikiANN), PosUDBOUN (UDTurkishBOUN), PosUDIMST (UDTurkishIMST). Table 7 presents the micro F1 scores. All Turkish models demonstrate high performance, with micro F1 scores exceeding $92$. The BERTurk model achieves the highest overall score of $93.67$, taking the leading position across all datasets. Consistent with the results observed in text classification tasks, TabiBERT maintains competitive performance levels.